OpenStudy (lilai3):

Please help me-Algebra 1! WILL FAN/REWARD MEDAL!!!!!!!!!!!!!!! Find all values of k so that the trinominal x^2 + kx -35 can be factored using integers.

OpenStudy (dan815):

all the factors of 35

OpenStudy (dan815):

find that and see all the possible sums

OpenStudy (lilai3):

all the possible factors of 35 are 1, 5, 7, and 35...

OpenStudy (dan815):

-35

OpenStudy (lilai3):

? i do not understand.

OpenStudy (whpalmer4):

Need the factors of -35, not 35

OpenStudy (dan815):

^

OpenStudy (lilai3):

Oh. Thanks. -1, -5, 7, 35...?

OpenStudy (lilai3):

But none of them works to make 1 in kx, though.

OpenStudy (anonymous):

I got k = +/- 34, +/- 2

OpenStudy (dan815):

basically -1*5*7=-35 or -1*1*35-35 so you can have -1 anywhere in between there so, -5+7 = -2 5-7=2 -35+1=-34 35-1 = 34

OpenStudy (anonymous):

as dan815 already explained :DD

OpenStudy (lilai3):

I do not understand the part where you said that -5 +7 = -2 so 5-7=2 5-7=-2, dan.

OpenStudy (dan815):

write out the possible integer factors without knowing k

OpenStudy (anonymous):

it came from factoring using AC method, when you have Ax^2 + Bx + C, and if this can be factors, B can be found two numbers that multiply to AC and add to B

OpenStudy (dan815):

(x-1)(x+35) = ? (x+1)(x-35) =? (x+7)(x-5)=? (x-7)(x+5)=?

OpenStudy (anonymous):

i in algebra 1 too! ;)

OpenStudy (dan815):

expand those brackets out

OpenStudy (dan815):

the constant comes from multiplying only the interger parts

OpenStudy (dan815):

the coefficient on the x^2 comes from multiplying only the x parts

OpenStudy (lilai3):